High Desert Potteryworks makes a variety of pottery products that it sells to retailers. The company uses a job-order costingRequired: 1. Compute the predetermined overhead rates used in the Molding Department and the Painting Department. 2. ComputeComplete this question by entering your answers in the tabs below. Req 1 Reg 2 Req 3A Req 3B Compute the total overhead costReq 1 Req 2 Req ЗА Req 3B What would be the total manufacturing cost recorded for Job 205? (Round Predetermined overhead ratReq 1 Req 2 Req 3A Req 3B If the job contained 35 units, what would be the unit product cost? (Round Predetermined overhead
Answer #1

1) Molding overhead rate = 222500/89000+2.60 = 5.10 per MH

Painting overhead rate = 470470/51700+4.60 = 13.70 per DLH

2) Total overhead cost = (320*5.1+13.70*133) = 3454

3) Total manufacturing cost

Molding Painting Total
Direct material 946 1280 2226
Direct labor 720 960 1680
Overhead 1632 1822 3454
Total 3298 4062 7360

4) Unit product cost = 7360/35 = 210.29